A 145 gram sample of a certain radioactive isotopes had a half-life of 500 years. If 4 half-lives go by, how much of the original isotope is left?


1
Expert's answer
2021-01-20T14:31:55-0500

One half-life: "\frac{145}{2}=72.5 \;g"

Two half-lives: "\frac{72.5}{2}=36.25 \;g"

Three half-lives: "\frac{36.25}{2}=18.125 \;g"

Four half-lives: "\frac{18.125}{2}=9.0625 \;g"

Answer: 9.0625 g